# Lepton Number Violation Higgs Decay at Muon Collider

### v1

@@ Abstract
In this paper, we consider the scalar singlet extension of type-I seesaw, where a scalar singlet $S$ and heavy neutral lepton $N$ are further introduced. The Majorana mass term of heavy neutral lepton is generated through the Yukawa interaction with the scalar singlet, which then induces the lepton number violation decays of SM Higgs $h$ and heavy Higgs $H$ via mixing of scalars. As a pathway to probe the origin of heavy neutral lepton mass, we investigate the lepton number violation Higgs decay signature at the TeV-scale muon collider. The dominant production channel of Higgs bosons at the TeV-scale muon collider is via vector boson fusion. So we perform a detailed analysis of the signal process $\mu^{+}\mu^{-}\to\nu_{\mu}\bar{\nu}_{\mu}h/H\to\nu_{\mu}\bar{\nu}_{\mu}NN$ followed by $N\to\mu^{\pm}jj$, where the two jets from $W$ boson decay are treated as one fat-jet $J$. With an integrated luminosity of $1(10)~{}\text{ab}^{-1}$, the 3 (10) TeV muon collider could discover the lepton number violation SM Higgs decay $h\to\mu^{\pm}\mu^{\pm}JJ$ signature for the Higgs mixing parameter $\sin\alpha>0.05(0.009)$. Meanwhile, a large parameter space can be detected by the lepton number violation heavy Higgs decay $H\to\mu^{\pm}\mu^{\pm}JJ$ signature for $m_{H}\lesssim 1(3)$ TeV and $\sin\alpha\gtrsim 0.03(0.005)$ at the 3 (10) TeV muon collider. Therefore, the lepton number violation SM and heavy Higgs decay signatures are both promising at the TeV scale muon collider.
@@ I Introduction
One natural pathway to explain the tiny neutrino mass is the seesaw mechanism Minkowski:1977sc ; Mohapatra:1979ia ; Schechter:1980gr ; Schechter:1981cv , which introduces heavy neutral lepton $N$. Due to the Majorana nature of $N$, one distinct signal is the lepton number violation signature at the collider, such as $pp\to W^{\pm(*)}\to\ell^{\pm}N\to\ell^{\pm}\ell^{\pm}jj$ Han:2006ip ; Deppisch:2015qwa ; Cai:2017mow . In the well-studied phenomenological type-I seesaw model, the lepton number violation signature is determined by the mixing parameter $V_{\ell N}$ between the light and heavy neutrinos Antusch:2016ejd ; Abdullahi:2022jlv . Currently, the LHC searches could exclude $|V_{\ell N}|^{2}\gtrsim 2\times 10^{-6}$ for prompt ATLAS:2019kpx ; CMS:2024xdq and $|V_{\ell N}|^{2}\gtrsim 3\times 10^{-7}$ for displaced signature ATLAS:2022atq ; CMS:2022fut ; CMS:2023jqi . However, the natural theoretical predicted squared mixing parameter is $|V_{\ell N}|^{2}\sim m_{\nu}/m_{N}<10^{-10}$ for $m_{N}>1$ GeV, which is even beyond the reach of future experiments Abdullahi:2022jlv .
@@
The Majorana mass term $m_{N}\bar{N}^{c}N$ is allowed by the SM gauge symmetry, so $m_{N}$ is usually assumed to be large enough to suppress the neutrino mass in the canonical seesaw. If we assign the lepton number $L=1$ to $N$, the Majorana mass term clearly breaks the lepton number symmetry by two units. Since $m_{N}\to 0$ will restore the lepton number symmetry, the bare Majorana mass term should be theoretically preferred to be small tHooft:1979rat , just such as in the inverse seesaw mechanism Mohapatra:1986bd . On the other hand, the origin of the Majorana mass term is not clarified in the minimal seesaw either. This term can be realized by extending the SM gauge symmetry, such as the left-right symmetrical model Pati:1974yy ; Mohapatra:1974gc ; Senjanovic:1975rk , the gauged $U(1)_{B-L}$ model Davidson:1978pm ; Marshak:1979fm ; Mohapatra:1980qe ; Liu:2021akf , the gauged $U(1)_{L_{\mu}-L_{\tau}}$ model Branco:1988ex ; He:1990pn ; Asai:2017ryy ; He:2024dwh ; Bi:2024pkk , and the general gauged $U(1)_{X}$ model Cox:2017eme ; Das:2017flq ; Das:2017deo ; Das:2022rbl . In such gauge symmetry extended models, heavy neutral lepton $N$ has additional interactions with the new gauge boson and new scalars. Searches of heavy neutral lepton via decays of new gauge bosons have been performed by LHC ATLAS:2023cjo ; CMS:2023ooo . For instance, the search for pair production of heavy neutral lepton via the process $pp\to Z^{\prime}\to NN$ with $N\to\ell^{\pm}jj$ by the CMS collaboration has excluded the parameter space with $m_{Z^{\prime}}\lesssim 4.4$ TeV and $m_{N}\lesssim 1.3$ TeV CMS:2023ooo .
@@
In this paper, we focus on the Higgs sector with an additional scalar singlet $S$ Shoemaker:2010fg ; Gao:2019tio , which usually appears in the $U(1)$ symmetric models. By assigning lepton number $L=-2$ to the new scalar $S$, the Majorana mass term then originates from the new Yukawa interaction $S\bar{N}^{c}N$, which is invariant under the $U(1)$ symmetry. After the spontaneous symmetry breaking, both the scalar singlet and the neutral component of SM Higgs doublet $\Phi$ obtain non-zero vacuum expectation values. The scalar interaction $(\Phi^{\dagger}\Phi)(S^{\dagger}S)$ induces the mixing between scalars. In the mass eigenstate, we denote $h$ to be the SM 125 GeV Higgs boson and $H$ to be the additional heavy Higgs boson. Including the mixing between scalars, both the SM $h$ and heavy Higgs boson $H$ couple to the heavy neutral lepton $N$, which could induce the lepton number violation Higgs decay as $h/H\to NN$ with $N\to\ell^{\pm}jj$ at colliders when kinematically allowed, i.e., $m_{N}<m_{h}/2$ or $m_{N}<m_{H}/2$ Maiezza:2015lza .
@@
The lepton number violation Higgs decay has been investigated at hadron Basso:2010yz ; Nemevsek:2016enw ; Accomando:2017qcs ; Fuks:2025jrn ; Liu:2025ldf and $e^{+}e^{-}$ colliders Gao:2021one ; Barducci:2020icf . In this paper, we study this signature at the recently proposed muon collider Delahaye:2019omf ; Accettura:2023ked ; InternationalMuonCollider:2024jyv ; InternationalMuonCollider:2025sys . Compared to the hadron collider, the muon collider has a cleaner environment. Meanwhile, the muon collider can achieve much higher center-of-mass energy than the $e^{+}e^{-}$ collider MuCoL:2024oxj , which also enables the reach of the heavy Higgs boson. Different from searches for heavy neutral lepton in canonical seesaw Mekala:2023diu ; Kwok:2023dck ; Li:2023tbx ; Li:2023lkl ; Cao:2024rzb , this signature is not suppressed by the mixing parameter $V_{\ell N}$. Meanwhile, instead of the Higgs-strahlung process at the Higgs factory Gao:2021one , the dominant production channel of Higgs bosons at the TeV scale muon collider is via vector boson fusion Han:2020pif ; Andreetto:2024rra , which provides a new independent pathway to probe the intrinsic of Higgs bosons. So the full signal process investigated in this paper is $\mu^{+}\mu^{-}\to\nu_{\mu}\bar{\nu}_{\mu}h/H\to\nu_{\mu}\bar{\nu}_{\mu}NN$ followed by $N\to\ell^{\pm}jj$.
@@ II The Model
The studied model in this paper is the type-I seesaw with an additional scalar singlet $S$. This singlet $S$ usually appears in $U(1)$ extension of seesaw models. To explain the smallness of neutrino mass, the seesaw mechanism works with heavy neutral lepton $N$ Minkowski:1977sc ; Mohapatra:1979ia ; Schechter:1980gr ; Schechter:1981cv , which carries lepton number $+1$. The scalar singlet $S$ with lepton number $-2$ not only breaks the $U(1)$ symmetry spontaneously, but also generates the Majorana mass of heavy neutral lepton via the Yukawa interaction $S\bar{N}^{c}N$. Hence, the mass of heavy neutral lepton $m_{N}$ is strongly associated with the scale of $U(1)$ symmetry, which we attribute to the TeV scale for collider searches. The local $U(1)$ scenario is considered in this paper to avoid the massless Goldstone boson. For different $U(1)$ symmetry, it can be distinguished from the interaction of new gauge boson due to different charge assignment of fermions KA:2023dyz , while the scalar sectors are similar in the $U(1)$ models.
@@
$V(\Phi,S)=m_{1}^{2}H\Phi^{\dagger}\Phi+m_{2}^{2}S^{\dagger}S+\lambda_{1}(\Phi^{\dagger}\Phi)^{2}+\lambda_{2}(S^{\dagger}S)^{2}+\lambda_{3}(\Phi^{\dagger}\Phi)(S^{\dagger}S),$ (1)
@@
The mixing between scalars modifies the coupling of SM Higgs by a factor of $\cos\alpha$, which then alerts the Higgs signal strengths. Using the precision Higgs measurements data ATLAS:2022vkf ; CMS:2022dwd , the global fit result of all channels requires $|\sin\alpha|<0.29$ Lane:2024vur ; Lewis:2024yvj . More stringent constraints come from the direct searches for the heavy Higgs in channels of $H\to ZZ$ CMS:2018amk ; ATLAS:2020tlo , $H\to WW$ CMS:2019bnu ; ATLAS:2020fry , and $H\to hh$ CMS:2021yci ; CMS:2022kdx ; ATLAS:2021ifb ; ATLAS:2022xzm . The combined analysis of all channels shows that $|\sin\alpha|\lesssim 0.2$ is allowed by direct searches Feuerstake:2024uxs . Meanwhile, the NLO calculations of $W$ boson mass set the most strict constraints for $m_{H}$ around the TeV scale, which could exclude $|\sin\alpha|\gtrsim 0.16$ Lopez-Val:2014jva ; Robens:2016xkb ; Feuerstake:2024uxs .
@@
Without fine-tuning, the typical value of the mixing parameter is $|V_{\ell N}|^{2}\sim m_{\nu}/m_{N}<10^{-12}$ for heavy neutral lepton above the electroweak scale. Such a small mixing parameter is far beyond the reach of current experimental limits ATLAS:2019kpx ; CMS:2024xdq .
@@ III Decay Properties
Theoretical predictions of BR$(h\to NN)$ are shown in Figure 1. With a typical value of $\sin\alpha\sim\mathcal{O}(0.1)$ under current experimental limits, the predicted branching ratio of $h\to NN$ is at the order of 0.01. As we fixed $v_{s}=1000$ GeV in this paper, decreasing $m_{N}$ will lead to smaller Yukawa coupling $y_{N}$, hence a smaller branching ratio. Meanwhile, the heavier $N$ has a smaller phase space of $h\to NN$, which results in the maximum branching ratio around $m_{N}=40$ GeV. It is clear in Equation (11) that the partial decay width $\Gamma(h\to NN)$ is proportional to $\sin^{2}\alpha$. When $\sin\alpha\gtrsim 0.05$, the predicted BR$(h\to NN)$ is larger than 0.001 for $m_{N}\sim 10-50$ GeV.
@@
When the heavy neutral lepton is lighter than the $W$ boson, the three body decays mediated by off-shell $W$ and $Z$ bosons are the dominant decay modes. Analytical expressions for the three body partial decay widths can be found in Ref. Atre:2009rg ; Liao:2017jiz . In this paper, we use Madgraph5_aMC@NLO Alwall:2014hca to evaluate the numerical results of branching ratios, which is shown in Figure 3. The three body decay mode $N\to\ell^{\pm}q\bar{q}^{\prime}$ has a branching ratio of about 0.5 when $m_{N}<m_{W}$. Meanwhile, the two body decay mode $N\to\ell^{\pm}W^{\mp}$ is larger than 0.5 when $m_{N}>m_{W}$. Hence, the fully visible decay mode $N\to\ell^{\pm}jj$ is always a dominant channel for $m_{N}>5$ GeV.

### v2

@@ Abstract
In this paper, we consider the scalar singlet extension of type-I seesaw, where a scalar singlet $S$ and a heavy neutral lepton $N$ are further introduced. The Majorana mass term of heavy neutral lepton is generated through the Yukawa interaction with the scalar singlet, which then induces the lepton number violation decays of SM Higgs $h$ and heavy Higgs $H$ via mixing of scalars. As a pathway to probe the origin of heavy neutral lepton mass, we investigate the lepton number violation Higgs decay signature at the TeV-scale muon collider. The dominant production channel of Higgs bosons at the TeV-scale muon collider is via vector boson fusion. So we perform a detailed analysis of the signal process $\mu^{+}\mu^{-}\to\nu_{\mu}\bar{\nu}_{\mu}h/H\to\nu_{\mu}\bar{\nu}_{\mu}NN$ followed by $N\to\mu^{\pm}jj$, where the two jets from $W$ boson decay are treated as one fat-jet $J$. With an integrated luminosity of $1(10)~\text{ab}^{-1}$, the 3 (10) TeV muon collider could discover the lepton number violation SM Higgs decay $h\to\mu^{\pm}\mu^{\pm}JJ$ signature for the Higgs mixing parameter $\sin\alpha>0.05(0.009)$. Meanwhile, a large parameter space can be detected by the lepton number violation heavy Higgs decay $H\to\mu^{\pm}\mu^{\pm}JJ$ signature for $m_{H}\lesssim 1(3)$ TeV and $\sin\alpha\gtrsim 0.03(0.005)$ at the 3 (10) TeV muon collider. Therefore, the lepton number violation SM and heavy Higgs decay signatures are both promising at the TeV scale muon collider.
@@ I Introduction
One natural pathway to explain the tiny neutrino mass is the seesaw mechanism Minkowski:1977sc ; Mohapatra:1979ia ; Schechter:1980gr ; Schechter:1981cv , which introduces heavy neutral lepton $N$. Due to the Majorana nature of $N$, one distinct signal is the lepton number violation signature at the collider delAguila:2005ssc ; Banerjee:2015gca ; Drewes:2019byd ; Mekala:2022cmm ; Antusch:2022ceb ; Mikulenko:2023ezx ; Mekala:2023kzo , such as $pp\to W^{\pm(*)}\to\ell^{\pm}N\to\ell^{\pm}\ell^{\pm}jj$ Han:2006ip ; Deppisch:2015qwa ; Cai:2017mow . In the well-studied phenomenological type-I seesaw model, the lepton number violation signature is determined by the mixing parameter $V_{\ell N}$ between the light and heavy neutrinos Antusch:2016ejd ; Abdullahi:2022jlv , which is mostly sensitive to the mass region with $m_{N}\lesssim 100$ GeV. Above the electroweak scale, $V_{\ell N}$ is less constrained as $|V_{\ell N}|^{2}\gtrsim 10^{-3}$ being disfavored via indirect constraints Antusch:2015mia . Currently, the LHC searches could exclude $|V_{\ell N}|^{2}\gtrsim 2\times 10^{-6}$ at $m_{N}=20$ GeV for prompt ATLAS:2019kpx ; CMS:2024xdq and $|V_{\ell N}|^{2}\gtrsim 3\times 10^{-7}$ at $m_{N}=12$ GeV for displaced signature ATLAS:2022atq ; CMS:2022fut ; CMS:2023jqi . In the future, CEPC and FCC-ee could probe $|V_{\ell N}|^{2}\gtrsim 10^{-10}$ with $m_{N}\sim$50 GeV Abdullahi:2022jlv , and the muon collider could test $|V_{\ell N}|\gtrsim 10^{-6}$ for TeV scale $m_{N}$ Mekala:2023diu ; Kwok:2023dck ; Li:2023tbx . However, the natural theoretical predicted squared mixing parameter is $|V_{\ell N}|^{2}\sim m_{\nu}/m_{N}<10^{-10}$ for $m_{N}>1$ GeV, which is even beyond the reach of future experiments Abdullahi:2022jlv .
@@
The Majorana mass term $m_{N}\bar{N}^{c}N$ is allowed by the SM gauge symmetry, so $m_{N}$ is usually assumed to be large enough to suppress the neutrino mass in the canonical seesaw. If we assign the lepton number $L=1$ to $N$, the Majorana mass term clearly breaks the lepton number symmetry by two units. Since $m_{N}\to 0$ will restore the lepton number symmetry, the bare Majorana mass term should be theoretically preferred to be small tHooft:1979rat , just such as in the inverse seesaw mechanism Mohapatra:1986bd . On the other hand, the origin of the Majorana mass term is not clarified in the minimal seesaw either. This term can be realized by extending the SM gauge symmetry, such as the left-right symmetrical model Pati:1974yy ; Mohapatra:1974gc ; Senjanovic:1975rk , the gauged $U(1)_{B-L}$ model Davidson:1978pm ; Marshak:1979fm ; Mohapatra:1980qe ; Liu:2021akf , the gauged $U(1)_{L_{\mu}-L_{\tau}}$ model Branco:1988ex ; He:1990pn ; Asai:2017ryy ; He:2024dwh ; Bi:2024pkk , and the general gauged $U(1)_{X}$ model Cox:2017eme ; Das:2017flq ; Das:2017deo ; Das:2022rbl ; Arun:2022ecj . In such gauge symmetry extended models, heavy neutral lepton $N$ has additional interactions with the new gauge boson and new scalars. Searches of heavy neutral lepton via decays of new gauge bosons have been performed by LHC ATLAS:2023cjo ; CMS:2023ooo . For instance, the search for pair production of heavy neutral lepton via the process $pp\to Z^{\prime}\to NN$ with $N\to\ell^{\pm}jj$ by the CMS collaboration has excluded the parameter space with $m_{Z^{\prime}}\lesssim 4.4$ TeV and $m_{N}\lesssim 1.3$ TeV CMS:2023ooo . In the future, the HL-LHC could probe the parameter space with $m_{Z^{\prime}}\lesssim 6$ TeV and $m_{N}\lesssim 3$ TeV Arun:2022ecj .
@@
In this paper, we focus on the Higgs sector with an additional scalar singlet $S$ Shoemaker:2010fg ; Gao:2019tio , which usually appears in the $U(1)$ symmetric models. By assigning lepton number L=2L=-2 to the new scalar $S$, the Majorana mass term then originates from the new Yukawa interaction $S\bar{N}^{c}N$, which is invariant under the $U(1)$ symmetry. After the spontaneous symmetry breaking, both the scalar singlet and the neutral component of SM Higgs doublet $\Phi$ obtain non-zero vacuum expectation values. The scalar interaction $(\Phi^{\dagger}\Phi)(S^{\dagger}S)$ induces the mixing between scalars. In the mass eigenstate, we denote $h$ to be the SM 125 GeV Higgs boson and $H$ to be the additional heavy Higgs boson. Including the mixing between scalars, both the SM $h$ and heavy Higgs boson $H$ couple to the heavy neutral lepton $N$, which could induce the lepton number violation Higgs decay as $h/H\to NN$ with $N\to\ell^{\pm}jj$ at colliders when kinematically allowed, i.e., $m_{N}<m_{h}/2$ or $m_{N}<m_{H}/2$ Maiezza:2015lza .
@@
The lepton number violation Higgs decay has been investigated at hadron colliders Basso:2010yz ; Nemevsek:2016enw ; Accomando:2017qcs ; Fuks:2025jrn ; Liu:2025ldf , which is sensitive to the parameter space with $m_{H}\lesssim 150$ GeV Nemevsek:2016enw . Meanwhile, the future $e^{+}e^{-}$ colliders could probe the mixing parameter between scalars $\sin\alpha$ to be lower than $10^{-2}$Gao:2021one ; Barducci:2020icf . In this paper, we study this signature at the recently proposed muon collider Delahaye:2019omf ; Accettura:2023ked ; InternationalMuonCollider:2024jyv ; InternationalMuonCollider:2025sys . Compared to the hadron collider, the muon collider has a cleaner environment. Meanwhile, the muon collider can achieve much higher center-of-mass energy than the $e^{+}e^{-}$ collider MuCoL:2024oxj , which also enables the reach of the heavy Higgs boson. Different from searches for heavy neutral lepton in canonical seesaw Mekala:2023diu ; Kwok:2023dck ; Li:2023tbx ; Li:2023lkl ; Cao:2024rzb , this signature is not suppressed by the mixing parameter $V_{\ell N}$. Meanwhile, instead of the Higgs-strahlung process at the Higgs factory Gao:2021one , the dominant production channel of Higgs bosons at the TeV scale muon collider is via vector boson fusion Han:2020pif ; Andreetto:2024rra , which provides a new independent pathway to probe the intrinsic of Higgs bosons. So the full signal process investigated in this paper is $\mu^{+}\mu^{-}\to\nu_{\mu}\bar{\nu}_{\mu}h/H\to\nu_{\mu}\bar{\nu}_{\mu}NN$ followed by $N\to\ell^{\pm}jj$.
Besides the neutral Higgs bosons $h/H$, there could be other origin of the lepton number violation signature at the muon collider, such as the new gauge boson $Z^{\prime}$ He:2024dwh and the charged Higgs boson $H^{\pm}$ Batra:2023ssq . The common same sign dilepton signature $\ell^{\pm}\ell^{\pm}4j$ could be induced as $\mu^{+}\mu^{-}\to\nu_{\mu}\bar{\nu}_{\mu}h/H\to\nu_{\mu}\bar{\nu}_{\mu}NN\to\ell^{\pm}\ell^{\pm}4j+\nu_{\mu}\bar{\nu}_{\mu}$ via the neutral Higgs bosons $h/H$, $\mu^{+}\mu^{-}\to Z^{\prime*}\to NN\to\ell^{\pm}\ell^{\pm}4j$ via the gauge boson $Z^{\prime}$, or $\mu^{+}\mu^{-}\to NN\to\ell^{\pm}\ell^{\pm}4j$ via the $t$-channel exchange of charged Higgs boson $H^{\pm}$. These origins can be distinguished via additional signatures. For instance, the new gauge boson could mediate the new signature $\mu^{+}\mu^{-}\to Z^{\prime}\gamma\to NN\gamma\to\ell^{\pm}\ell^{\pm}4j+\gamma$ He:2024dwh . Meanwhile the charged Higgs boson can be pair produced at muon collider, which leads to the new signature as $\mu^{+}\mu^{-}\to H^{+}H^{-}\to\ell^{+}N+\ell^{-}N\to\ell^{+}\ell^{-}\ell^{\pm}\ell^{\pm}4j$ Batra:2023ssq .
@@ II The Model
The studied model in this paper is the type-I seesaw with an additional scalar singlet $S$. This singlet $S$ usually appears in $U(1)$ extension of seesaw models. To explain the smallness of neutrino mass, the seesaw mechanism works with heavy neutral lepton $N$ Minkowski:1977sc ; Mohapatra:1979ia ; Schechter:1980gr ; Schechter:1981cv , which carries lepton number +1+1. The scalar singlet $S$ with lepton number 2-2 not only breaks the $U(1)$ symmetry spontaneously, but also generates the Majorana mass of heavy neutral lepton via the Yukawa interaction $S\bar{N}^{c}N$. Hence, the mass of heavy neutral lepton $m_{N}$ is strongly associated with the scale of $U(1)$ symmetry, which we attribute to the TeV scale for collider searches. The local $U(1)$ scenario is considered in this paper to avoid the massless Goldstone boson Chikashige:1980qk ; Pilaftsis:1993af . For different $U(1)$ symmetry, it can be distinguished from the interaction of new gauge boson due to different charge assignment of fermions KA:2023dyz , while the scalar sectors are similar in the $U(1)$ models.
@@
$V(\Phi,S)=m_{1}^{2}\Phi^{\dagger}\Phi+m_{2}^{2}S^{\dagger}S+\lambda_{1}(\Phi^{\dagger}\Phi)^{2}+\lambda_{2}(S^{\dagger}S)^{2}+\lambda_{3}(\Phi^{\dagger}\Phi)(S^{\dagger}S),$ (1)
@@
The mixing between scalars modifies the coupling of SM Higgs by a factor of $\cos\alpha$, which then alters the Higgs signal strengths. In this paper, we consider the constraints from current LHC searches. Using the precision Higgs measurements data ATLAS:2022vkf ; CMS:2022dwd , the global fit result of all channels requires $|\sin\alpha|<0.29$ Lane:2024vur ; Lewis:2024yvj . More stringent constraints come from the direct searches for the heavy Higgs in channels of $H\to ZZ$ CMS:2018amk ; ATLAS:2020tlo , $H\to WW$ CMS:2019bnu ; ATLAS:2020fry , and $H\to hh$ CMS:2021yci ; CMS:2022kdx ; ATLAS:2021ifb ; ATLAS:2022xzm . The combined analysis of all channels shows that $|\sin\alpha|\lesssim 0.2$ is allowed by direct searches Feuerstake:2024uxs , where the explicit constraint on $\sin\alpha$ depends on the mass of heavy Higgs. Meanwhile, the NLO calculations of $W$ boson mass set the most strict constraints for $m_{H}$ around the TeV scale, which could exclude $|\sin\alpha|\gtrsim 0.16$ Lopez-Val:2014jva ; Robens:2016xkb ; Feuerstake:2024uxs .
@@
Without fine-tuning, the typical value of the mixing parameter is $|V_{\ell N}|^{2}\sim m_{\nu}/m_{N}<10^{-12}$ for heavy neutral lepton above the electroweak scale. In this paper, we fix the mixing parameter as $|V_{\ell N}|^{2}=m_{\nu}/m_{N}=10^{-12}\times(50~\text{GeV})/m_{N}$ with $m_{\nu}=0.05$ eV in the following studies. For instance, $m_{N}=50(500)$ GeV corresponds to $|V_{\ell N}|^{2}=10^{-12}(10^{-13})$. Such small mixing parameters are far beyond the reach of current experimental limits ATLAS:2019kpx ; CMS:2024xdq .
@@ III Decay Properties
[…]
@@
When the heavy neutral lepton is lighter than the $W$ boson, the three body decays mediated by off-shell $W$ and $Z$ bosons are the dominant decay modes. Analytical expressions for the three body partial decay widths can be found in Ref. Atre:2009rg ; Liao:2017jiz . In this paper, we use Madgraph5_aMC@NLO Alwall:2014hca to evaluate the numerical results of branching ratios, which is shown in Figure 3. The three body decay mode $N\to\ell^{\pm}q\bar{q}^{\prime}$ has a branching ratio of about 0.5 when $m_{N}<m_{W}$. Meanwhile, the two body decay mode $N\to\ell^{\pm}W^{\mp}$ is larger than 0.5 when $m_{N}>m_{W}$. Hence, the fully visible signature $N\to\ell^{\pm}jj$ is always a dominant channel for $m_{N}>5$ GeV.